# Introspective Planning: Guiding Language-Enabled Agents to Refine Their Own Uncertainty

### v2

@@ Abstract
Large language models (LLMs) exhibit advanced reasoning skills, enabling robots to comprehend natural language instructions and strategically plan high-level actions through proper grounding. However, LLM hallucination may result in robots confidently executing plans that are misaligned with user goals or, in extreme cases, unsafe. Additionally, inherent ambiguity in natural language instructions can induce task uncertainty, particularly in situations where multiple valid options exist. To address this issue, LLMs must identify such uncertainty and proactively seek clarification. This paper explores the concept of introspective planning as a systematic method for guiding LLMs in forming uncertainty–aware plans for robotic task execution without the need for fine-tuning. We investigate uncertainty quantification in task-level robot planning and demonstrate that introspection significantly improves both success rates and safety compared to state-of-the-art LLM-based planning approaches. Furthermore, we assess the effectiveness of introspective planning in conjunction with conformal prediction, revealing that this combination yields tighter confidence bounds, thereby maintaining statistical success guarantees with fewer superfluous user clarification queries.
Large Language Model Reasoning, Large Language Model in Robotics, Conformal Prediction, Uncertainty Quantification
@@ 1 Introduction
Figure 1: Illustration of the introspective planning pipeline. Knowledge base construction: the LLM is prompted to generate knowledge entries based on human-provided instructions and the correct options. Deployment: Upon receiving an instruction, the LLM formulates possible next steps and consults the knowledge base to retrieve the most relevant examples, which are later used as the prompt for prediction.
@@
Large Language Models (LLMs), when pre-trained on internet-scale text corpora, demonstrate emergent capabilities that extend far beyond mere text comprehension and generation as their scale increases (Brown et al., 2020; Wei et al., 2022a). Through prompting (Gao et al., 2021) and in-context learning (Min et al., 2022), these models have shown remarkable adaptability ranging from answering complex questions and solving mathematical problems to generating computer code and engaging in sophisticated reasoning processes during inference (Srivastava et al., 2023). Robots interacting with humans can leverage the capabilities of LLMs to interpret task instructions in natural language, employ common sense reasoning to understand their environment, and devise high-level action plans grounded in the capabilities and affordances of the robot (Ahn et al., 2022; Liang et al., 2023).
@@
The reliability of LLM outputs has direct implications downstream robotics tasks. Language models are prone to hallucinations (Ji et al., 2023a), which cause models to generate plans that are at odds with common-sense knowledge, not executable by the robot, or incompatible with the environment constraints (Zeng et al., 2023). For example, if a human user asks a robot to bake some bread in a kitchen containing an oven and varied cookware, the robot’s LLM may generate a decision to use a plastic tray without considering the risk of it melting. Furthermore, possible ambiguities in the user’s request can also introduce uncertainty into the LLM’s reasoning and planning (Tamkin et al., 2023). In our example, while multiple containers are suitable for the stated task, biases inherited from training data may tilt action generation towards certain options. Therefore, the robot needs to calibrate its uncertainty quantification and seek further communication with users when ambiguities are identified.
Human beings assess their internal values and knowledge of their own abilities to guide domain-level reasoning processes: this is referred to as introspective reasoning (Leake, 2012). In this paper, we observe that LLMs can leverage an analogous reasoning scheme to better assess underlying uncertainty when generating plans. In addition, drawing inspiration from the concept of self-explanation (Dunlosky et al., 2013), we proposed a novel LLM planning technique. This technique emulates the introspective and self-explanatory reasoning of humans, suggesting that prompting LLMs with ground truth options and encouraging them to consider the reasons behind these options can lead to the creation of more accurate and reliable plans. This methodology significantly departs from traditional retrieval methods in automatic reasoning (Zhang et al., 2023) by specifically addressing uncertainty and safety in task-level planning. It not only facilitates more nuanced decision-making by the LLMs but also aligns closely with human-like introspective reasoning processes, thereby enhancing the interpretability and reliability of these decisions. We also integrated introspective planning with conformal prediction approaches (Angelopoulos & Bates, 2021; Angelopoulos et al., 2022) to provide a statistical guarantee on task completion. Our evaluation shows that introspective planning empirically outperforms the state of the art (Ren et al., 2023) across metrics and asks for clarification less to ensure a target success rate.
Statement of contributions. (1) We demonstrate the benefits of systematically guiding LLMs to account for their own uncertainty in plan generation. Introspective reasoning enhances LLM planning quality, surpasses state-of-the-art techniques, and improves interpretability via generated explanatory feedback. (2) We augment mobile manipulation datasets with safety-relevant scenarios and propose new metrics reflecting a planner’s specification compliance and degree of conservativeness, complementary to previously measured task completion. We conduct thorough evaluations that highlight the robustness of introspective planning in diverse scenarios. (3) We show that, when used with conformal prediction, introspective planning makes fewer user queries to maintain its statistical guarantee, and readily handles ambiguous task statements in its multi-label setting.
@@ 2 Introspective Planning
Problem Formulation. Similar to (Ren et al., 2023), we cast LLM-based planning as multiple-choice question answering (MCQA). Given a task statement $d_{i}$ and the observation $o_{i}$, the LLM planner first generates a set of candidate plans $\mathcal{C}_{i}$, each assigned a unique letter label from $\mathcal{Y}\coloneqq\{\textrm{`}A\textrm{'},\textrm{`}B\textrm{'},\textrm{`}C\textrm{'},\ldots\}$ ’ , … }. The planner then predicts $\hat{y}_{i}\in\mathcal{Y}$, aiming to match the unknown true user intent $z_{i}$. For example, consider the stated task $d_{i}$ring me that soda” with the observation $o_{i}$ that a banana, a pack of chips, and a can of Coke are placed on the counter. The LLM planner will first generate three options $\mathcal{C}_{i}$ of bringing each item to the user, and predict the label $\hat{y}_{i}$ corresponding to the Coke .
@@
Planning with Knowledge Retrieval. At inference time, the planner selects the most pertinent reasoning examples from the knowledge base $\mathcal{K}$ to aid the LLM’s reasoning. Given a test instance ${x}_{\text{test}}=({d}_{\text{test}},{o}_{\text{test}})$, we compute the cosine similarity between the text embedding of ${d}_{\text{test}}$ and all keys of $\mathcal{K}$. As shown in Figure 1, we retrieve the most relevant knowledge corresponding to the $m$ most similar embeddings as prompt and leverage the in-context learning capabilities of the LLM to generate possible plans and reason about their feasibility. To select the desired robot plan $\hat{y}_{\text{test}}$ with generated reasoning, we can use two distinctive prediction methods. 1) Direct Prediction: We ask the LLM to output the best option $\hat{y}_{\text{test}}$ along with all possible plans and explanations. 2) Conformal Prediction: Instead of directly predicting $\hat{y}_{\text{test}}$, we construct a set of valid candidate plans ${\hat{\mathcal{G}}}_{\text{test}}\subseteq\mathcal{C}_{\text{test}}$ by querying LLM’s confidence $\hat{f}(y_{i}|{x}_{\text{test}},\mathcal{C}_{\text{test}},{k}_{\text{test}})$ for each label $y_{i}\in\mathcal{Y}$ given the prompt constructed by knowledge retrieval process and generated reasoning. The robot will request human for help if multiple valid options are included in the ${\hat{\mathcal{G}}}_{\text{test}}$. In the following section, we demonstrate how to incorporate introspective planning with conformal prediction.
@@
$\mathcal{U}=\{(x_{1},\mathcal{G}_{1},),\ldots,(x_{N},\mathcal{G}_{N},)\}$$\mathcal{K}\leftarrow\{\}$Initialize Knowledge Base Dictionary each train example $x_{i}=(d_{i},o_{i})$$\mathcal{C}_{i}\leftarrow\text{GenerateChoices}(x_{i})$$k_{i}\leftarrow\text{GenerateKnowledge}(x_{i},\mathcal{C}_{i},\mathcal{G}_{i})$$e_{i}\leftarrow\text{Embed}(d_{i})$$\mathcal{K}[e_{i}]\leftarrow\{x_{i},\mathcal{C}_{i},k_{i},\mathcal{G}_{i}\}$
@@
Algorithm 2 Introspective Conformal Prediction
${x}_{\text{test}}=({d}_{\text{test}},{o}_{\text{test}})$, $\mathcal{K}$, $m$, $\hat{q}$${e}_{\text{test}}\leftarrow\text{Embed}({d}_{\text{test}})$$\mathcal{K}_{\text{similar}}\leftarrow\text{FindSimilar}({e}_{\text{test}},\mathcal{K}.\text{keys()},m)$$\mathcal{C}_{\text{test}},{k}_{\text{test}}\leftarrow\text{Plan}({x}_{\text{test}},\mathcal{K}_{\text{similar}})$${\hat{\mathcal{G}}}_{\text{test}}\leftarrow\text{ConformalPrediction}({x}_{\text{test}},\mathcal{C}_{\text{test}},{k}_{\text{test}},\hat{q})$$|\hat{\mathcal{G}}_{\text{test}}|==1$${y}_{\text{test}}\leftarrow\hat{\mathcal{G}}_{\text{test}}$equest for further instructions.
[…]
@@ 3 Introspective Conformal Prediction
Figure 2: Demonstration of using conformal prediction with Introspective Planning. After generating multiple options, we query the LLM for the explanation by introspective planning and then ask the model to predict the most correct option. Based on the likelihood scores of true intents from a calibration dataset, conformal prediction finds the quantile value $\hat{q}$ (0.85), and includes any options scoring above $\geq 1-\hat{q}$ in the prediction set for each test scenario. This method guarantees the correct answer is included among the options, at a confidence level specified by the user.
@@
This approach ensures the coverage criterion specified in Equation 1, providing a statistically justified guarantee for the comprehensiveness of the prediction set.
As the marginal guarantee in Equation 1 depends on both the calibration set and the test set, every time we have a new test instance $z_{test}$, we would ideally sample a new calibration set to maintain the same level of statistical assurance, which could be too resource-intensive. However, we can choose $N$ large enough to control the fluctuations in coverage by analyzing its distribution. The distribution of coverage has an analytic form as follows (Vovk, 2012):
@@
where $l=\lfloor(N+1)\hat{\epsilon}\rfloor$, $\text{Beta}_{N+1-l,l}^{-1}(\delta)$ denotes the inverse CDF (quantile) level of $\delta$ in a Beta distribution with parameters $N+1-l$ and $l$, and $\hat{\epsilon}$ is the threshold used for calibration. Additionally, prior research by Sadinle (2019) (Sadinle et al., 2019) demonstrates that conformal prediction minimizes the prediction set size, suggesting that robots employing this method require the least human intervention while attaining desired success rates.
Prior work KnowNo (Ren et al., 2023) utilizes a similar conformal prediction approach for planning. In this work, we significantly enhance this framework by incorporating introspective planning rationale $k_{i}$ to improve the likelihood function’s effectiveness. This adjustment optimizes the distribution of nonconformity scores, leading to a tighter concentration around smaller values. Such refinement leads to tighter bounds, improving the framework’s reliability and reducing its conservativeness, as will be demonstrated in subsequent sections. We summarize the procedure of introspective conformal prediction in Algorithm 2.
Extending to Multi-Label Prediction. In many cases, the user’s task statements can truly be ambiguous, and this means that there’ll be multiple distinct candidate plans that comply with what the user asked for in different ways, even though the user may only be satisfied with one of them. Therefore the robot needs to be able to reason about how certain it is that the task was specified ambiguously.
@@
4 Evaluation Method
Figure 3: Qualitative results on Safe Mobile Manipulation. We compared our approach with KnowNo (Ren et al., 2023). Both approaches use conformal prediction with 85% target success rate. Our approach generates the explanation first through introspective planning and then predict the valid options by conformal prediction while KnowNo directly predicts the valid options through conformal prediction. We observed significant improvement in generating more precise prediction sets using introspective planning.
4.1 Evaluation Metrics
Traditional measures such as success rate and help rate inadequately capture the nuances of LLM behavior, particularly in scenarios requiring precision and accuracy. We introduce additional metrics to capture prediction’s overconfidence in ambiguous tasks and imprecision when the LLM mixes valid and invalid options in their planning.
  • Success Rate (SR) measures how often the language model’s predictions match the user’s intent, calculated as the percentage of cases where the predicted actions include the correct intent.
  • Help Rate (HR) is the percentage of cases where the prediction set encompasses more than one option, such that robots will require further human instructions.
  • Exact Set Rate (ESR) measures the frequency of the LLM’s predictions perfectly aligning with all valid actions inferred from instructions. It’s crucial for evaluating the model’s ability to generate precise responses.
  • Non-compliant Contamination Rate (NCR) calculates the proportion of prediction sets containing options that deviate from the given instructions, thereby measuring the LLM’s ability to follow directions accurately.
  • Unsafe Contamination Rate (UCR) measures the frequency at which the prediction sets include potentially unsafe options, assessing the model’s ability to prioritize safety in responses.
4.2 Evaluation Baselines
We benchmarked our proposed introspective planning against various prompt-based methods to gauge its effectiveness in LLM reasoning. The Prompt Set (Ren et al., 2023) instruct the LLM to directly output the prediction through few-shot in-context learning. Prompt Set+CoT (Wei et al., 2022b) applies a Chain of Thought (CoT) process to simultaneously produce explanations and predictions. Retrieval-Q-CoT (Zhang et al., 2023) utilizes CoT to generate reasoning in a training dataset and retrieves the most relevant prompt during inference. Auto-CoT (Zhang et al., 2023) automates prompts selecting process by using clustering to ensure a broad representation of diverse scenarios.
To further illustrate the effectiveness of introspective planning in enhancing conformal prediction, we compared our method with KnownNo (Ren et al., 2023), which integrates conformation prediction with Prompt Set. Additionally, we employed conformal calibration for Retrieval-Q-CoT (Zhang et al., 2023) serving as an extra baseline due to its use of retrieval augmentation for reasoning. All calibration processes used the same dataset with 400 instances, and the same hyper-parameter $\delta=0.01$ consistent with original implementation in KnowNo.

### v3

# Introspective Planning: Aligning Robots’ Uncertainty with Inherent Task Ambiguity

@@ Abstract
Large language models (LLMs) exhibit advanced reasoning skills, enabling robots to comprehend natural language instructions and strategically plan high-level actions through proper grounding. However, LLM hallucination may result in robots confidently executing plans that are misaligned with user goals or, in extreme cases, unsafe. Additionally, inherent ambiguity in natural language instructions can induce task uncertainty, particularly in situations where multiple valid options exist. To address this issue, LLMs must identify such uncertainty and proactively seek clarification. This paper explores the concept of introspective planning as a systematic method for guiding LLMs in forming uncertainty–aware plans for robotic task execution without the need for fine-tuning. We investigate uncertainty quantification in task-level robot planning and demonstrate that introspection significantly improves both success rates and safety compared to state-of-the-art LLM-based planning approaches. Furthermore, we assess the effectiveness of introspective planning in conjunction with conformal prediction, revealing that this combination yields tighter confidence bounds, thereby maintaining statistical success guarantees with fewer superfluous user clarification queries. Code is available at https://github.com/kevinliang888/IntroPlan.
@@ 1 Introduction
Large Language Models (LLMs), when pre-trained on internet-scale text corpora, demonstrate emergent capabilities that extend far beyond mere text comprehension and generation as their scale increases [5, 40]. Through prompting [9] and in-context learning [26], these models have shown remarkable adaptability ranging from answering complex questions and solving mathematical problems to generating computer code and engaging in sophisticated reasoning processes during inference [36]. Robots interacting with humans can leverage the capabilities of LLMs to interpret task instructions in natural language, employ common sense reasoning to understand their environment, and devise high-level action plans grounded in the capabilities and affordances of the robot [1, 21].
Figure 1: Illustration of the introspective planning pipeline. Knowledge base construction: The LLM generates knowledge entries based on human-provided instructions and correct options. Deployment: Upon receiving an instruction, the LLM formulates possible next steps, consults the knowledge base to retrieve the most relevant examples, and uses them as prompts for prediction.
@@
The reliability of LLM outputs has direct implications for downstream robotics tasks. Language models are prone to hallucinations [13], which cause models to generate plans that are at odds with common-sense knowledge, not executable by the robot, or incompatible with the environment constraints [48]. For example, if a human user asks a robot to bake some bread in a kitchen containing an oven and varied cookware, the robot’s LLM may generate a decision to use a plastic tray without considering the risk of it melting. Furthermore, possible ambiguities in the user’s request can also introduce uncertainty into the LLM’s reasoning and planning [37]. In our example, while multiple containers are suitable for the stated task, biases inherited from training data may tilt action generation towards certain options. Therefore, the robot needs to calibrate its uncertainty quantification and seek further communication with users when ambiguities are identified.
Human beings assess their internal values and knowledge of their own abilities to guide domain-level reasoning processes: this is referred to as introspective reasoning [19]. In this paper, we observe that LLMs can leverage an analogous reasoning scheme to better assess underlying uncertainty when generating plans. We propose a novel method for constructing a knowledge base that utilizes LLMs to generate human-aligned introspective reasoning examples with minimal human input. During inference, this human-aligned knowledge guides LLMs to produce more reliable and interpretable plans. Unlike traditional Retrieval-Augmented Generation (RAG) approaches [10, 20, 29, 43, 51], which utilize open-source, off-the-shelf knowledge bases to enhance text generation, our approach retrieves few-shot introspective reasoning examples from the knowledge base. This enables LLMs to explicitly reason about uncertainties and formulate plans in a structured format. Additionally, our method augments previous automatic reasoning approaches [49] by integrating human feedback into the reasoning generation process. We have observed that introspective planning, when integrated with conformal prediction [2, 3], refines the LLM’s uncertainty and provides a tighter statistical guarantee.
Statement of contributions.
To the best of our knowledge, this is the first work to integrate retrieval-augmented planning with conformal prediction, refining language agents’ uncertainty and reducing user queries while maintaining statistical guarantees. Key contributions are summarized as follows:
  • We propose a novel introspective planning scheme that prompts language-enabled agents to proactively assess their own confidence regarding task compliance and safety for multiple candidate plans, with a guaranteed probability that the agent will either execute the actions desired by the user or ask an appropriate follow-up question to disambiguate the user’s intent.
  • We introduce a new, weakly supervised offline knowledge base construction method that guides the LLM to generate human-aligned introspective reasoning examples as post-hoc rationalizations of human-selected safe-and-compliant plans.
  • We create a new Safe Mobile Manipulation benchmark, which augments previous mobile manipulation datasets with safety-critical scenarios and introduces new metrics to evaluate a planner’s specification compliance, safety, and degree of conservativeness.
@@ 2 Introspective Planning
Problem Formulation. Similar to [31], we cast LLM-based planning as multiple-choice question answering (MCQA). Given a task statement $d_{i}$ and the observation $o_{i}$, the LLM planner first generates a set of candidate plans $\mathcal{C}_{i}$, each assigned a unique letter label from $\mathcal{Y}\coloneqq\{\textrm{`}A\textrm{'},\textrm{`}B\textrm{'},\textrm{`}C\textrm{'},\ldots\}$ ’ , … }. The planner then predicts $\hat{y}_{i}\in\mathcal{Y}$, aiming to match the unknown true user intent $z_{i}$. For example, consider the stated task $d_{i}$ring me that soda" with the observation $o_{i}$ that a banana, a pack of chips, and a can of Coke are placed on the counter. The LLM planner will first generate three options $\mathcal{C}_{i}$ of bringing each item to the user, and predict the label $\hat{y}_{i}$ corresponding to the Coke.
@@
Planning with Knowledge Retrieval. At inference time, the planner selects the most pertinent reasoning examples from the knowledge base $\mathcal{K}$ to aid the LLM’s reasoning. Given a test instance ${x}_{\text{test}}=({d}_{\text{test}},{o}_{\text{test}})$, we compute the cosine similarity between the text embedding of ${d}_{\text{test}}$ and all keys of $\mathcal{K}$. As shown in Figure 1, we retrieve the most relevant knowledge corresponding to the $m$ most similar embeddings as prompt and leverage the in-context learning capabilities of the LLM to generate possible plans and reason about their feasibility. To select the desired robot plan $\hat{y}_{\text{test}}$ with generated reasoning, we can use two distinctive prediction methods: (1) Direct Prediction: We ask the LLM to output the best option $\hat{y}_{\text{test}}$ along with all possible plans and explanations. (2) Conformal Prediction: Instead of directly predicting $\hat{y}_{\text{test}}$, we construct a set of valid candidate plans ${\hat{\mathcal{G}}}_{\text{test}}\subseteq\mathcal{C}_{\text{test}}$ by querying LLM’s confidence $\hat{f}(y_{i}|{x}_{\text{test}},\mathcal{C}_{\text{test}},{k}_{\text{test}})$ for each label $y_{i}\in\mathcal{Y}$ given the prompt constructed by knowledge retrieval process and generated reasoning. The robot will request human for help if multiple valid options are included in the ${\hat{\mathcal{G}}}_{\text{test}}$. In the following section, we demonstrate how to incorporate introspective planning with conformal prediction.
@@
1:$\mathcal{U}=\{(x_{1},\mathcal{G}_{1}),\ldots,(x_{N},\mathcal{G}_{N})\}$
2:$\mathcal{K}\leftarrow\{\}$triangleright Knowledge Base Initialization
3:for each train example $x_{i}=(d_{i},o_{i})$ do
4:     $\mathcal{C}_{i}\leftarrow\text{GenerateChoices}(x_{i})$
5:     $k_{i}\leftarrow\text{GenerateKnowledge}(x_{i},\mathcal{C}_{i},\mathcal{G}_{i})$
6:     $e_{i}\leftarrow\text{Embed}(d_{i})$
7:     $\mathcal{K}[e_{i}]\leftarrow\{x_{i},\mathcal{C}_{i},k_{i},\mathcal{G}_{i}\}$
8:end for
Algorithm 2 Introspective Conformal Prediction
1:${x}_{\text{test}}=({d}_{\text{test}},{o}_{\text{test}})$, $\mathcal{K}$, $m$, $\hat{q}$
2:${e}_{\text{test}}\leftarrow\text{Embed}({d}_{\text{test}})$
@@
4:$\mathcal{C}_{\text{test}},{k}_{\text{test}}\leftarrow\text{Plan}({x}_{\text{test}},\mathcal{K}_{\text{similar}})$
5:${\hat{\mathcal{G}}}_{\text{test}}\leftarrow\text{ConformalPred}({x}_{\text{test}},\mathcal{C}_{\text{test}},{k}_{\text{test}},\hat{q})$
6:if $|\hat{\mathcal{G}}_{\text{test}}|==1$ then ${y}_{\text{test}}\leftarrow\hat{\mathcal{G}}_{\text{test}}$
7:else  Request further instructions
8:end if
@@ 3 Introspective Conformal Prediction
Figure 2: Demonstration of using conformal prediction with Introspective Planning. After generating multiple options, we query the LLM for the explanation by introspective planning and then ask the model to predict the most correct option. Based on the likelihood scores of true intents from a calibration dataset, conformal prediction finds the quantile value $\hat{q}$ (0.85), and includes any options scoring above $\geq 1-\hat{q}=0.15$ in the prediction set for each test scenario. This method guarantees the correct answer is included among the options, at a confidence level specified by the user.
@@
This approach ensures the coverage criterion specified in Equation 1, providing a statistically justified guarantee for the comprehensiveness of the prediction set. The proof is shown in Appendix E.
As the marginal guarantee in Equation 1 depends on both the calibration set and the test set, every time we have a new test instance $z_{test}$, we would ideally sample a new calibration set to maintain the same level of statistical assurance, which could be too resource-intensive. However, we can choose $N$ large enough to control the fluctuations in coverage by analyzing its distribution. The distribution of coverage has an analytic form as follows [38]:
@@
where $l=\lfloor(N+1)\hat{\epsilon}\rfloor$, $\text{Beta}_{N+1-l,l}^{-1}(\delta)$ denotes the inverse CDF (quantile) level of $\delta$ in a Beta distribution with parameters $N+1-l$ and $l$, and $\hat{\epsilon}$ is the threshold used for calibration. Additionally, prior research by Sadinle (2019) [32] demonstrates that conformal prediction minimizes the prediction set size, suggesting that robots employing this method require the least human intervention while attaining desired success rates.
Prior work KnowNo [31] utilizes a similar conformal prediction approach for planning. In this work, we significantly enhance this framework by incorporating introspective planning rationale $k_{i}$ to improve the likelihood function’s effectiveness. This adjustment optimizes the distribution of nonconformity scores, leading to a tighter concentration around smaller values. Such refinement leads to tighter bounds, improving the framework’s reliability and reducing its conservativeness, as will be demonstrated in subsequent sections. We summarize the procedure of introspective conformal prediction in Algorithm 2.
4 Evaluation
4.1 Evaluation Method
Asking for help is not enough. Previous work uses success rate and help rate as metrics, but these do not fully capture a planner’s performance. For example, if the instruction is to bring the soda and the robot’s prediction set includes Coke, Sprite, and Apple, the robot will ask for help but ask the wrong question due to an irrelevant option. Neither the success rate nor the help rate captures this issue because ’success’ is only defined as the prediction set containing the user intent. Additionally, help rate can sometimes be misleading. A low help rate does not necessarily indicate good performance, as an effective predictor should ideally seek help whenever instructions are ambiguous.
To address these, we categorized errors into three types: (1) The robot is uncertain, but the task is unambiguous. (2) The robot is certain but wrong. (3) The robot is uncertain, and the task is ambiguous, but it asks the wrong question. Based on this analysis, we proposed new metrics to capture these errors. Exact set rate and non-compliant contamination rate effectively measure the error type (3). Overask rate captures the error type (1) while the overstep rate captures the error type (2). Additionally, we propose Unsafe Contamination Rate (UCR) and Unsafe Rate to measure the robot’s performance in prioritizing safety, which previous metrics do not account for.
Figure 3: Qualitative results on Safe Mobile Manipulation. We compared our approach with KnowNo [31], both using conformal prediction with an 85% target success rate. Our method generates explanations via introspective planning before applying conformal prediction, whereas KnowNo directly predicts valid options using conformal prediction. We observed that KnowNo over-step in the left case and over-ask in the right case while IntroPlan generates more precise prediction sets.
Metrics. Beyond the success rate and help rate, we introduce additional metrics to more comprehensively evaluate the performance of our planner.
  • Success Rate (SR): How often the language model’s predictions match the user’s intent, calculated as the percentage of cases where the predicted actions include the correct intent.
  • Help Rate (HR): Fraction of cases where the prediction set encompasses more than one option, such that robots will require further human instructions, $\text{HR}=N_{\text{ask}}/N$.
  • Exact Set Rate (ESR): Frequency of the LLM’s predictions perfectly aligning with all valid actions inferred from instructions. It evaluates the model’s ability to generate precise responses.
  • Non-compliant Contamination Rate (NCR): Proportion of prediction sets containing options that deviate from the given instructions, measuring the LLM’s ability to follow instructions accurately and ask the right questions to clarify uncertainty.
  • Unsafe Contamination Rate (UCR): Frequency at which the prediction sets include potentially unsafe options, assessing the model’s ability to prioritize safety in responses.
  • Overask Rate: Fraction of instances when the planner is uncertain while the task is unambiguous.
@@ 4.1 Evaluation Method
[…]
  • Unsafe Rate: Frequency at which planner is certain to execute unsafe action.
Baselines. We benchmarked our proposed introspective planning against various prompt-based methods to gauge its effectiveness in LLM reasoning. The Prompt Set [31] instruct the LLM to directly output the prediction through few-shot in-context learning. Prompt Set+CoT [41] applies a Chain of Thought (CoT) process to simultaneously produce explanations and predictions. Retrieval-Q-CoT [49] utilizes CoT to generate reasoning in a training dataset and retrieves the most relevant prompt during inference. Auto-CoT [49] automates prompts selecting process by using clustering to ensure a broad representation of diverse scenarios.
To further illustrate the effectiveness of introspective planning in enhancing conformal prediction, we compared our method with KnownNo [31], which integrates conformation prediction with Prompt Set. Additionally, we employed conformal calibration for Retrieval-Q-CoT [49] serving as an extra baseline due to its use of retrieval augmentation for reasoning. All calibration processes used the same dataset with 400 instances, and the same hyper-parameter $\delta=0.01$ consistent with original implementation in KnowNo.